Service Observing for DTime UT Date 07102016

Mountian: David

Tucson: Jenny, Shane, Andrew, Greg, Xianyu.

mods1b&r.20160710.0011- imflats in ugri filters for Wen-fai.

18:04 LUCI1 darks

18:07 MODS1 grating flats

18:17 MODS1R fits not flushing, David checking mods1.rc. Not at DOS prompt but in a weird state. David restarting IC process.

MODS1R file 0024 is test exposure.

MODS1R 0021 is missing header info.

18:41 manually repeating MODS1R flats. Previous flats had vflat = 10 instead of 5.

18:44 Flats still saturated, rerunning script. OK now.

19:02 Running LBC biases. Skies are clear so we are going to attempt LBC flats.

19:28 mods1r file 0035 has a bad row (y=1376). We saw this a couple of images prior as well.

19:30 closed dome calibrations finished. Opening. LUCI1 dark 0064 is still being taken, don\x92t use.

19:35 Sunset

19:41 Slewing to twilight field.

19:53 Testing LBC flat levels for B and R, still saturating.

20:00 Still saturating.

20:01 First good flats for B.

20:02 32k B, 40k R. Pixels dropping out on R\x85 025932

20:06 Scaling factor 2.

20:07 Scaling factor 4.

20:10 Scaling factor 7.

20:17 Scaling factor 17. Twilights finished

20:17 Slewing to Bootes co-pointing ob

20:26 Running WRS.

20:32 12 degree twilight

21:00 WRS did not give us sensible results, the telescope was still defocused after it ran. Running dofpia.

21:09 We were running the wrong script, loaded co-point.ob and starting again.

21:10 18 degree twilight

21:16 Slewing to science target

21:25 ~0.7\x94 on Blue and Red Bootes images

21:32 ~ 0\x94.9 on .

21:42 Seeing appears stable. ~ 0\x94.8 on the fourth set of red and blue images. 0\x94.8 on DIMM.

21:45 Starting the Bootes scripts for the second and final time. Last images from first pass are still around 0\x94.8 - 0\x94.9. Surprisingly seeing on the blue images is a little better than that on the red, perhaps dofpia did not do a great job focusing the red side.

22:03 Half way through the second pass of the script, still ~ 0\x94.8 on both sides.

22:08 LBC obs finished, reconfiguring for LUCI for UAO

23:11 UOA configured, shell set. Preset sent.

23:40 UAO has not been successful, Greg and Xianyu are perplexed. They are handing back the telescope. Seeing may or may not be good now, DIMM is sporadic but we are at high elevation. David is testing at lower elevation.

23:45 Seeing is ~1.3 so interrupt is not going ahead.

23:52 Slewing to Olga\x92s LUCI target q2048.

00:02 We need to home the AGW, we can see its thermal signature at the bottom of the LUCI field. Acquisition was started, aiming to acquire at 1086, 1024.

00:09 Xianyu is leaving for the night.

00:10 Calculated offset is 1\x94.67, 8\x94.39

00:12 Cancelling preset, homing AGW, sending preset again, sending previously calculated offset and skipping to through slit image.

00:14 The thermal emission remains, we need to change guide star

00:23 Having double checked with Olga she is OK with us going ahead with the same guide star.

00:24 Science exposure started.

00:55 Seeing has been variable, adding a few extra exposures.

01:15 Slewing to telluric HIP102074.

01:27 Starting telluric exposure

01:41 Slewing to q2158

02:00 Starting science exposures. Acquisition offset was -0.57, -5.01. Target was acquired on 1086, 1023 and then shifted 15\x94 up.

02:14 1\x94.4 - 1\x94.6 on the DIMM, 1\x94.2 - 1\x94.4 on the guider.

02:37 As the script moved from zJ to HK the wavelength failed to set correctly. We set it manually, reset the script, skipped to the first HK exposure and started it from there. Note that this means the offset of -7 in Y will have been applied twice, but as we started 15\x94 higher than normal I don\x92t see that as being a problem.

02:51 The DIT and NDIT were not correct for HK and sky lines are saturated (File 0103). The script has 1x450s, the README says it should be 4x100s. Editing script and starting again. 900s lost.

03:30 Slewing to telluric HIP109452

03:34 Acquired on 1079.5, 1023.5. Offset was 0\x94.12, -5\x94.38.

03:13 Starting telluric exposure.

03:15 Spectrum is saturated, adjusting exposure time.

03:28 Returning to q2158

03:30 18 degree twilight

03:43 Starting science

04:12 Skipping one of the 450s zJ exposures.

04:17 12 degree twilight. Science completed, closing up.

04:26 Starting arcs, flats. Saving darks for the afternoon. Had a saturated flat, HALO1 was on instead of HALO2.

04:55 Also running MODS1 cals to test the IC.

05:14 Sunrise

HKspec 1\x94 flats 1.93: 6.5s Halo2. Files 153-162

Arcs for above, neon, argon, xenon. Files 163-168

zJspec 1\x94 flats 1.17, 25s Halo2. Files 183-192

Arcs for above, neon, argon, zenon. Files 193-200

0\x94.5 G210 H 1.65, 2.9s Halo1 (could use 2.7s next time). Files 233-242

Arcs for above, neon, argon, zenon. Files 243 - 250

H band imaging flats. 3.8s in Halo3. Files 263 - 272

-- %USERSIG{AndrewCardwell - 2016-07-10}%

Comments


 
Topic revision: r1 - 10 Jul 2016, AndrewCardwell
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback